OpenStudy (anonymous):

steve and tom sold 104 raffle tickets. steve sold 17 more than twice as many raffle tickets as tom. how many raffle tickets did each boy sell

OpenStudy (anonymous):

\[steve=2tom+17\]\[steve+tom=104\]\[2tom+17=104-tom\]\[tom=87/3\]

OpenStudy (anonymous):

let no of tickets sold by Tom = x no of tickets sold by Steve = 2x + 17 x + 2x + 17 = 104 3x = 104 - 17 x = 87/3 x = 29 so Tom sold 29 tickets Steve sold 75 tickets